# Automated Structural-level Alignment of Multi-view TLS and ALS Point Clouds in Forestry

### v3

@@ 3.2.2 Non-uniform ground terrain
icp [44] 2.324 0.627 25.488 0.88 10.104 0.979
cpd [45] 0.0374 5.3638 29.8273 4.5839 10.0638 0.7630
gmm [46] 0.136 0.0398 18.311 4.583 10.063 0.763
svr [47] 1.724 3.781 23.909 5.388 9.166 0.478
gmmtree [48] 1.590 3.952 28.412 4.6770 8.501 0.618
filterReg [49] 0.512 6.764 30.691 4.451 12.503 0.903
bcpd [50] 0.2248 5.062 21.402 3.2667 12.6039 2.4279
ours 0.0224 0.0002 0.0059 0.0156 0.0149 0.0180
Table 3 illustrates the superior performance of our method in registering heterogeneous forest scenes compared to standard methods. The table highlights that while some methods known for their effectiveness in urban or man-made object scenarios perform well, they encounter challenges in forestry applications. Our approach addresses these challenges by leveraging incremental alignments based on the levels of structural complexity in the point-cloud pairs.
@@ 3.3 Pair-wise TLS to ALS co-registration
Validation of our pair-wise structural complexity approach was also performed for TLS-to-ALS co-registration. Two datasets were used: the NM dataset and the Fort Stewart dataset. The NM dataset includes co-located ALS scans along with a total of 46 TLS scans covering the region. In Fort Stewart, GA, there were 42 plot-scale TLS point clouds with regional-scale ALS data. For ALS cropping, cropped scans of 50x50 m centered around the approximate GPS location of each TLS scan center were used (used solely for ALS region cropping). TLS plots were registered against ALS, with all six alignment parameters randomly initialized simultaneously from a uniform distribution centered at the GPS position of the center scan, deviating by $\pm$15m for translation and $\pm$45 for rotation. We used three levels of structural complexity $K=3$ to group the TLS scans and two levels $K=2$ for ALS. A representative example demonstrating the regional coverage from the Fort Stewart dataset and the result of applying our method to co-register TLS scans with ALS is shown in Fig.7. In the figure, the yellow-colored point cloud represents the regional ALS scan, while the differently colored blobs depict individual TLS point cloud scans with a radius of approximately 50 meters each. The middle column provides a zoomed-in view highlighting the regions of overlap between scans, while the right-most column shows a side-by-side comparison of the co-registered ALS (top) and TLS (bottom) scans at the plot scale. Despite the different spatial resolutions—ALS appearing sparser—our approach effectively aligns both point clouds, as evident from the zoomed-in comparison in the right-most column at the forest plot scale. In terms of computational performance, it took approximately 3 minutes to co-register each pair of scans, totaling 95 minutes to co-register the entire set of 31 pairs of scans in the Fort Stewart dataset.
@@
Quantitative evaluation of TLS-to-ALS co-registration performance was conducted using registration RMSE relative to ALS orientation as ground truth, as described in Section 3.1 (Note that target-based methods for obtaining ground truth is impractical in TLS and ALS co-registration because of a 2 year time disparity between the ALS and the TLS collection). We evaluated performance in both the NM and the Ft. Stewart datasets, where for each pair we generate a total of 100 random initializations. Fig. 8 illustrates the distribution of initializations (in blue) and the error estimates (in orange) from our proposed approach. The vertical axis represents error frequency, while the horizontal axis shows each alignment parameter out of the six. Fig.8.a,c summarize the rotation and translation results for the Ft. Stewart dataset, and Fig. 8.b,d for the NM dataset, respectively. In all cases, the orange plots corresponding to estimation error are tightly concentrated around zero with minimal variability. RMSE values displayed in the horizontal labels for each parameter demonstrate that our approach achieves precise registration relative to ground truth. In both datasets, our method achieved an RMSE of less than $0.6^{\circ}$ for rotation and $0.07$m for translation. Slightly better performance in the Ft. Stewart dataset can be attributed to its higher ALS spatial resolution of approximately 15 points per square meter compared to 10 points per square meter in the NM dataset, partly due to some areas being surveyed twice in the former dataset. It is noteworthy that registration errors in ALS-to-TLS are larger compared to pair-wise TLS-to-TLS co-registration, which we attribute to TLS’s higher spatial resolution compared to ALS in general. For structural complexity settings, we employed $K=3$ levels for grouping TLS scans and $K=2$ for ALS. This allowed our approach to effectively match 3D points associated with lower complexity (e.g., ground) between ALS and TLS, while leaving out TLS points corresponding to tree trunks/branches, which are sparse or absent in ALS due to its challenges in returning points from the mid-story, as observed in the upper-right part of Fig. 7.
@@ 4 Discussion
Empirical evidence shows that when we set the number of structural complexity groups to $K=3$, it tends to group 3D points mainly into ground, tree trunks/branches, and foliage. A representative example illustrated in Fig.9.(a-c) corroborates this. Here, 3D points color-coded in yellow correspond to the ground surface associated with the group of lowest complexity, 3D points color-coded in beige represent the foliage at the highest complexity, while those color-coded in red correspond to tree trunks/branches at mid-structural complexity. This grouping is characteristic of our approach in TLS when $K=3$. We point out in Fig.9.(b), that points in groups of complexity 1 and 3 mainly consist of ground and foliage, similar to the point clouds captured by ALS, characterized by a sparse resolution in the mid-story as seen in the ALS side view in the top-right of Fig.7. This finding was exploited by our approach when co-registering TLS to ALS. Specifically, our assignment problem in Eq.(3) matched the highest complexity in TLS to that of ALS. In the case of $K=2$ shown in Fig.9(d-f), we observed that the mixture of vMF distribution groups 3D points mainly into two groups: ground and above-ground vegetation.
@@
(b) Complexity 1 and 3: $K=3$
(c) Complexity 1,2 and 3: $K=3$
(d) Complexity 1: $K=2$
(e) Complexity 2: $K=2$
(f) Complexity 1 and 2: $K=2$
Figure 9: Representative 3D points are grouped by structural complexity into different sets of $K$-groups. For $K=2$, our approach automatically groups 3D points into ground and above-ground vegetation. Similarly, with $K=3$, points are classified into ground, foliage, and tree trunks/branches/small shrubs, effectively partitioning the underlying forest components.
@@
Comparisons between the results of the mixture of vMF distributions in Fig. 9 and the mixture of Gaussians for splitting point clouds and computing structural complexity are shown in Fig.10. Note that the $k$-means approach was not able to split the point cloud into the underlying forest composition. This provides evidence that the mixture of Gaussians does not offer a natural clustering into structural complexities. The uneven sampling of 3D points within the range of the LiDAR sensor complicates the formation of fair groups of structural complexities. The standard $k$-means algorithm, assuming a mixture of Gaussians, results in clusters that reduce variance evenly as they assume iid 3D point samples. Overall, we found that our structural complexity metrics and grouping approach provide a natural way to segment point clouds from forests by their underlying components (e.g., ground, tree trunks/branches, foliage) without requiring prior knowledge about their shape, as is typically assumed by standard segmentation approaches, nor the use of learning-based methods, which typically require many labeled training example point clouds, which can become impractical. One drawback of our method that we would like to point out is that the number of $K$-groups to which we categorize 3D points into structural complexity is generally unknown and can vary depending on the characteristics of the natural ecosystem. In our experiments so far, we used $K=2$ and $K=3$ throughout all cases, which resulted in good alignment performance overall. However, the number of groups is a tunnable parameter that has to be tested by trial and error. In the following, we present some additional co-registration experiments in ecosystems where there is no structure from trees and where a different number of structural complexity groups makes more sense.
@@
One drawback of our proposed method is that it takes approximately 5 minutes to complete a single TLS pair-wise co-registration. This may not be a problem when only a few pairs of point clouds are available, as is typically the case with existing standard forest surveying practices. However, scaling to thousands or even millions of point cloud pairs can become impractical with the current implementation. To this end, improvements to computational efficiency could be explored by taking advantage of parallelization strategies, which could significantly reduce the convergence time of our proposed approach.
@@
In terms of additional benefits, we would like to point out that our innovation enables the capability to produce integrated LiDAR maps automatically with the attractive flexibility of access to broad low-res ALS data to rapidly query ecosystem attribute estimation summaries and at the same time access to detailed co-registered high-resolution in-situ understory TLS data for more detailed refined inference of characterizations at the available sites. This flexibility offers the potential to exploit and demonstrate the advantages of either individual or integrated LiDAR sensing sources for ecosystem analysis while also opens up avenues to define the necessary conditions and best strategies for more efficient sensing in ecosystem monitoring. In summary, our co-registration approach is an enabling technology for large scale ecosystem monitoring through more advanced computational tools of data source integration while also offering new possibilities to more advanced analysis at scale. It also enables the ability to monitor sites with a more efficient framework to update ecosystem information (as in the pre/post Rx fire co-registation) and with this, the capability to track localized changes over time. We think our co-registration methodology can also help in establishing trade-offs between ALS, TLS and integrated ALS+TLS based monitoring comparisons and relative to standard forest plot surveying. This, measured trough comparisons by time efficiency, human labor requirements, associated costs, performance accuracy, scalability. In future work, we intend to explore a comprehensive investigation of these tradeoffs through the aforementioned metrics. These capabilities altogether can overall improve planning, management and more efficient decision-making of natural ecosystem lands.

### v4

@@ 3.2.2 Non-uniform ground terrain
icp [48] 1.587 0.822 23.488 3.38 10.516 0.491
cpd [49] 0.093 3.81 22.949 4.865 10.371 0.654
gmm [50] 0.153 0.127 21.921 4.263 10.611 0.524
svr [51] 1.81 3.250 25.481 5.162 10.731 0.761
gmmtree [52] 1.542 4.373 27.952 4.677 9.667 0.591
filterReg [53] 0.373 5.982 29.055 4.362 10.332 0.823
bcpd [54] 0.328 4.173 20.940 4.384 10.010 2.182
Manual co-Reg 0.21 0.27 0.052 0.033 0.036 0.002
(CloudCompare [44])
Trunk-based TFR [25] 1.62 2.13 3.1 0.281 0.32 0.051
Tree locations [47] 1.27 2.35 4.15 0.322 0.419 0.073
Canopy-based [34] 3.59 4.62 3.1 0.546 0.419 0.103
ForestAlign (ours) 0.224 0.261 0.0059 0.0356 0.0349 0.0018
The second row block in Table 3 includes results from direct manual co-registration (i.e., note that the method to obtain ground truth using cross-validation is different), as well as the results of the TFR method [25], the tree locations based from [47] and the canopy-based method from [34], the final block presents the performance of our approach. We would like to note that in this case, significant occlusion from trees and understory debris impacted the scans due to the vegetation density of the ecosystem. The TFR method [25] encountered challenges in finding reliable correspondences between trees using the DBH metric. It generally performed closer to the ground truth in areas with sparse understory vegetation and fewer trees but struggled in cases with dense vegetation, primarily because the prediction of tree metrics, such as DBH, becomes problematic with heavy occlusion. This is not the case of course, in forests with a sparse spatial distribution of trees as in the benchmark dataset in Table 2 where TFR performed generally well. The tree location method performed similarly to the TFR method. The reason is that both being trunk-based will benefit or degrade depending on the trunk segmentation and feature extraction estimation. The canopy-based method from [34] also faced challenges with co-registration for this dataset. TLS scans taken from distances of 5-15 meters view the scene from different perspectives, and under dense forest conditions, occlusion causes significant variation in canopy feature estimates depending on scan position. In contrast, the proposed approach performed comparably to the manual co-registration method and generally exhibited small errors relative to the ground truth. We attribute this performance to the fact that our method does not rely on an intermediate feature estimation stage, which can potentially lose the rich information contained in the full point cloud. This advantage is particularly important as the accurate estimation of tree features becomes increasingly challenging and more prone to error with higher vegetation density and scene complexity.
@@ 3.3 Pair-wise TLS to ALS co-registration
Validation of our pair-wise structural complexity approach was also performed for TLS-to-ALS co-registration. Two datasets were used: the NM dataset and the Fort Stewart dataset. The NM dataset includes co-located ALS scans along with a total of 46 TLS scans covering the region. In Fort Stewart, GA, there were 42 plot-scale TLS point clouds with regional-scale ALS data. For ALS cropping, cropped scans of 50x50 m centered around the approximate GPS location of each TLS scan center were used (used solely for ALS region cropping). TLS plots were registered against ALS, with all six alignment parameters randomly initialized simultaneously from a uniform distribution centered at the GPS position of the center scan, deviating by $\pm$15m for translation and $\pm$45 for rotation. We used three levels of structural complexity $K=3$ to group the TLS scans and two levels $K=2$ for ALS. A representative example demonstrating the regional coverage from the Fort Stewart dataset and the result of applying our method to co-register TLS scans with ALS is shown in Fig.8. In the figure, the yellow-colored point cloud represents the regional ALS scan, while the differently colored blobs depict individual TLS point cloud scans with a radius of approximately 50 meters each. The middle column provides a zoomed-in view highlighting the regions of overlap between scans, while the right-most column shows a side-by-side comparison of the co-registered ALS (top) and TLS (bottom) scans at the plot scale. Despite the different spatial resolutions, ALS appearing sparser, our approach effectively aligns both point clouds, as evident from the zoomed-in comparison in the right-most column at the forest plot scale. In terms of computational performance, it took approximately 3 minutes to co-register each pair of scans, totaling 95 minutes to co-register the entire set of 31 pairs of scans in the Fort Stewart dataset.
@@
Quantitative evaluation of TLS-to-ALS co-registration performance was conducted using registration RMSE relative to ALS orientation as ground truth, as described in Section 3.1 (Note that target-based methods for obtaining ground truth is impractical in TLS and ALS co-registration because of a 2 year time disparity between the ALS and the TLS collection). We evaluated performance in both the NM and the Ft. Stewart datasets, where for each pair we generate a total of 100 random initializations. Fig. 9 illustrates the distribution of initializations (in blue) and the error estimates (in orange) from our proposed approach. The vertical axis represents error frequency, while the horizontal axis shows each alignment parameter out of the six. Fig.9.a,c summarize the rotation and translation results for the Ft. Stewart dataset, and Fig. 9.b,d for the NM dataset, respectively. In all cases, the orange plots corresponding to estimation error are tightly concentrated around zero with minimal variability. RMSE values displayed in the horizontal labels for each parameter demonstrate that our approach achieves precise registration relative to ground truth. In both datasets, our method achieved an RMSE of less than $0.6^{\circ}$ for rotation and $0.075$m for translation. It is noteworthy that registration errors in TLS-to-ALS are slighlty larger compared to pair-wise TLS-to-TLS co-registration, which we attribute to TLS’s higher spatial resolution compared to ALS in general. For structural complexity settings, we employed $K=3$ levels for grouping TLS scans and $K=2$ for ALS. This allowed our approach to effectively match 3D points associated with lower complexity (e.g., ground) between ALS and TLS, while leaving out TLS points corresponding to tree trunks/branches, which are sparse or absent in ALS due to its challenges in returning points from the mid-story, as observed in the upper-right part of Fig. 8.
Tables 4-5 present a comparison between forest-based co-registration methods and the proposed approach. The methods compared include the trunk-based TFR method [25], the tree-locations method [47], and the canopy-based method [34] across the NM and Ft. Stewart datasets. From the results, the trunk-based TFR and tree-locations methods faced the greatest challenges. This difficulty arises primarily because ALS data, with its limited ability to penetrate the canopy, produces a sparse distribution of 3D points in the mid-story, offering limited information. Of the two methods, TFR appeared to perform slightly better, likely due to fewer challenges in estimating DBH compared to determining relative tree positions and correspondences. In contrast, the canopy-based method and the proposed approach demonstrated more favorable co-registration error. ALS data provides uniform spatial resolution across the canopy and is less affected by occlusion at this level, facilitating canopy feature extraction and reducing co-registration error. For the TLS-to-ALS case, the proposed approach exploited on this property by matching the high structural complexity of canopy points in the TLS scan with corresponding high-complexity points in the ALS data, primarily from the tree canopy. Additionally, it leveraged ground topography information to obtain a rough initial co-registration estimate, which was subsequently refined using matched points at higher complexities, primarily in the canopy. The superior performance of the proposed method can be attributed to this approach—utilizing both ground and canopy 3D points. Overall, the proposed method outperformed the tree-based approaches and, similar to the TLS-to-TLS case, compared favorably to manual co-registration in the NM dataset.
Table 4: Registration RMSE comparison in TLS-to-ALS in NM dataset
Orientation (Degrees) Position (meters)
$\theta_{\text{roll}}$ $\theta_{\text{pitch}}$ $\theta_{\text{yaw}}$ $t_{x}$ $t_{y}$ $t_{z}$
CloudCompare [44] 0.231 0.341 0.391 0.042 0.041 0.024
TFR [25] 1.567 2.374 3.432 0.612 0.482 0.674
Tree-locations [47] 1.865 2.912 3.914 0.528 0.51 0.614
Canopy-based [34] 1.067 0.91 0.57 0.268 0.317 0.395
ForestAlign 0.771 0.532 0.263 0.0632 0.065 0.033
Table 5: Registration RMSE comparison in TLS-to-ALS in Ft. Stewart dataset
Orientation (Degrees) Position (meters)
$\theta_{\text{roll}}$ $\theta_{\text{pitch}}$ $\theta_{\text{yaw}}$ $t_{x}$ $t_{y}$ $t_{z}$
TFR [25] 1.142 0.935 2.14 0.31 0.481 0.181
Tree-locations [47] 2.754 1.59 1.82 0.48 0.53 0.3
Canopy-based [34] 0.981 1.03 0.79 0.271 0.162 0.147
ForestAlign 0.622 0.561 0.321 0.063 0.078 0.011
@@ 4 Discussion
Qualitative empirical evidence shows that when we set the number of structural complexity groups to $K=3$, it tends to group 3D points mainly into ground, tree trunks/branches/shrubs, and foliage. A representative example illustrated in Fig.10.(a-e) corroborates this. Here, 3D points color-coded in yellow correspond to the ground surface associated with the group of lowest complexity, 3D points color-coded in beige represent the foliage at the highest complexity, while those color-coded in red correspond to tree trunks/branches at mid-structural complexity. This grouping is characteristic of our approach in TLS when $K=3$. We point out in Fig.10.(d), that points in groups of complexity 1 and 3 mainly consist of ground and foliage, similar to the point clouds captured by ALS, characterized by a sparse resolution in the mid-story as seen in the ALS side view in the top-right of Fig.8. This finding was exploited by our approach when co-registering TLS to ALS. Specifically, our assignment problem in Eq.(3) matched the lowest and highest complexity in TLS to those of ALS. In the case of $K=2$ shown in Fig.10(f-h), we observed that the mixture of vMF distribution groups 3D points mainly into two groups: ground and above-ground vegetation.
@@
(b) Complexity 2: $K=3$
(c) Complexity 3: $K=3$
(d) Complexity 1,3: $K=3$
(e) Complexity 1-3: $K=3$
(f) Complexity 1: $K=2$
(g) Complexity 2: $K=2$
(h) Complexity 1 and 2: $K=2$
Figure 10: Representative 3D points are grouped by structural complexity into different sets of $K$-groups. For $K=2$, our approach automatically groups 3D points into ground and above-ground vegetation. Similarly, with $K=3$, points are classified into ground, tree trunks/branches/small shrubs, and foliage effectively partitioning the underlying forest components.
@@
Comparisons between the results of the mixture of vMF distributions in Fig. 10 and the mixture of Gaussians for splitting point clouds and computing structural complexity are shown in Fig.11. Note that the $k$-means approach was not able to split the point cloud into the underlying forest composition. This provides evidence that the mixture of Gaussians does not offer a natural clustering into structural complexities. The uneven sampling of 3D points within the range of the LiDAR sensor complicates the formation of fair groups of structural complexities. The standard $k$-means algorithm, assuming a mixture of Gaussians, results in clusters that reduce variance evenly as they assume independent and identically distributed 3D point samples. Overall, we qualitatively found that our structural complexity metrics and grouping approach splits point clouds from forests by their underlying components (e.g., ground, tree trunks/branches, foliage).
One drawback of our method that we would like to point out is that the number of $K$-groups to which we categorize 3D points into structural complexity is generally unknown and can vary depending on the characteristics of the natural ecosystem. In our experiments so far, we used $K=2$ for ALS and $K=3$ for TLS throughout all cases, which resulted in good alignment performance overall. However, the number of groups $K$ is a parameter that has to be tuned for performance by trial and error. In the following, we present some additional co-registration experiments in ecosystems where there is no structure from trees and where a $K=2$ for TLS made more sense.
@@
In terms of robustness to plausible erroneous initial estimates, we would like to highlight that our method is local, meaning it was designed to co-register point clouds miss-aligned by deviations less than 45° in rotations and a few meters in translations. Under these conditions the directional ambiguity often encountered when co-registering two planes (e.g., grounds) using standard approaches, such as ICP, does not arise. In our case, no issues were observed during the first stage of co-registration, where the lowest complexity features (typically the ground) were used to obtain a rough estimate of the co-registration parameters. Moreover, our approach assumes natural forests exhibit some variability in topography. Even in the Eastern US test dataset at Ft. Stewart, Georgia, where the terrain is known to be practically flat, our approach successfully detected subtle variability in the ground to obtain an initial co-registration estimate. Finally, the incremental nature of our approach, which aggregates points sequentially as complexity increases has a built-in capability to recover from erroneous initial estimates. Thus, if the ground were flat and an erroneous estimate were to be made at the lower complexity level, the method can recover at the next complexity level, where additional points above the ground can help to further constrain all six degrees of freedom characterizing the alignment between point clouds.
In the case where the ground is completely flat and lacks any vertical variability (e.g., uniform ground and grass), and without a good initial estimate of the co-registration parameters, our method would be able to align up to the $x,y$ dimensions due to the lack of constraints on all six degrees of freedom. However, we would like to contend that such scenarios are highly unlikely in real-world forest settings. In dense forest scenarios, the robustness of our approach depends on the structural diversity of the landscape. If sufficient topographic variability exists, the method can leverage these subtleties to constrain all six degrees of freedom and achieve a good initial alignment, which can then be further refined, similar to the Ecuador case, as long as the search space for correspondences remains within a close neighborhood. This also applies to dense foliage, regardless of understory structure. However, in the unlikely case of a completely flat ground with no variability to constrain all six degrees of freedom and dense, uniformly spaced vegetation (i.e., uniform tree spacing and foliage distribution), our approach may encounter difficulties. This challenge arises because the initial stages may not provide a good initial alignment, and subsequent, higher-complexity stages may fail to refine or lock the co-registration parameters at a global maximum. Instead, the dense foliage or uniform tree spacing may create numerous local minima in the optimization function, resulting in likely incorrect alignments. In intermediate scenarios, where natural forests exhibit some variability in topography and/or in the tree-stand direction, no issues were observed with co-registration.
In summary, tree-based methods perform well in some cases; however, their reliance on sparse tree-based features derived from the point cloud, rather than the entire point cloud, makes their accuracy highly dependent on the pre-processing stage used to estimate these tree-based features. This line of approaches sacrifice the level of spatial detail and feature richness of the point cloud, limiting the overall accuracy of the co-registration method. Moreover, increased occlusion in the scene generally leads to larger errors in tree feature predictions, which, in turn, introduce errors in the co-registration optimization. This is problematic in both TLS-to-TLS and TLS-to-ALS scenarios, in the former case as occlusion increasingly affects the level of information as a function of range, and in the later because of the LiDAR’s limited ability to capture beyond the canopy. In contrast, our method leverages the full information available in the point cloud, similar in spirit to ICP but better suited for scenes with varying levels of structural complexity. Furthermore, our approach is capable of co-registering natural scenes even in the absence of trees (e.g., in prairies), as long as there is at least some subtle elevation variability in the topography. This is because our method can use ground variability to find rough alignments, unlike tree-based methods, which typically do not exploit ground features for co-registration. Moreover, our approach can exploit different levels of structural complexity to achieve co-registration in both TLS-to-TLS and TLS-to-ALS scenarios. In the TLS-to-TLS case, it implicitly leverages the advantages of trunk-based methods by initially focusing on under-canopy features to obtain a rough alignment. In the TLS-to-ALS case, it starts by focusing on ground variability features to achieve a rough alignment, and then refines the alignment using features around the canopy level. This is an advantage that tree-feature based methods do not provide, and rather, one has to use different co-registration methods on a case by case basis. Overall, our approach outperformed forest-semantic-based co-registration methods and achieved comparable results to manual co-registration and target-based methods, with the added advantage of being fully automatic and target-free.
@@
One additional drawback of the proposed method is that it takes approximately 5 minutes to complete a single TLS pair-wise co-registration. In the case of TLS-to-ALS co-registration it was approximately 3 minutes. This may not represent a problem when only co-registering a few pairs of point clouds, as is typically the case with existing standard forest surveying practices. However, scaling to thousands or even millions of point cloud pairs can become impractical with the current implementation. To this end, improvements to computational efficiency could be explored by taking advantage of parallelization strategies, which could significantly reduce the convergence time of ForestAlign.
@@
In terms of additional benefits, we would like to point out that our innovation enables the capability to produce integrated LiDAR maps automatically with the attractive flexibility of access to broad low-res ALS data to rapidly query ecosystem attribute estimation summaries and at the same time access to detailed co-registered high-resolution in-situ understory TLS data for more detailed refined inference of characterizations at the available sites. This flexibility offers the potential to exploit and demonstrate the advantages of either individual or integrated LiDAR sensing sources for ecosystem analysis while also opens up avenues to define the necessary conditions and best strategies for more efficient sensing in ecosystem monitoring. The proposed co-registration approach is an enabling technology for large scale ecosystem monitoring through more advanced computational tools of data source integration while also offering new possibilities to more advanced analysis at scale. It also enables the ability to monitor sites with a more efficient framework to update ecosystem information (as in the pre/post Rx fire co-registation) and with this, the capability to track localized changes over time. We think our co-registration methodology can also help in establishing trade-offs between ALS, TLS and integrated ALS+TLS based monitoring comparisons and relative to standard forest plot surveying. This, measured trough comparisons by time efficiency, human labor requirements, associated costs, performance accuracy, scalability. In future work, we intend to explore a comprehensive investigation of these tradeoffs through the aforementioned metrics. These capabilities altogether can overall improve planning, management and more efficient decision-making of natural ecosystem lands.